# MVCNet: Multi-View Contrastive Network for Motor Imagery Classification

### v3

@@ Abstract
Electroencephalography (EEG)-based brain-computer interfaces (BCIs) enable neural interaction by decoding brain activity for external communication. Motor imagery (MI) decoding has received significant attention due to its intuitive mechanism. However, most existing models rely on single-stream architectures and overlook the multi-view nature of EEG signals, leading to limited performance and generalization. We propose a multi-view contrastive network (MVCNet), a dual-branch architecture that parallelly integrates CNN and Transformer models to capture both local spatial-temporal features and global temporal dependencies. To enhance the informativeness of training data, MVCNet incorporates a unified augmentation pipeline across time, frequency, and spatial domains. Two contrastive modules are further introduced: a cross-view contrastive module that enforces consistency of original and augmented views, and a cross-model contrastive module that aligns features extracted from both branches. Final representations are fused and jointly optimized by contrastive and classification losses. Experiments on five public MI datasets across three scenarios demonstrate that MVCNet consistently outperforms seven state-of-the-art MI decoding networks, highlighting its effectiveness and generalization ability. MVCNet provides a robust solution for MI decoding by integrating multi-view information and dual-branch modeling, contributing to the development of more reliable BCI systems.
@@ I Introduction
A further limitation of current MI decoding networks lies in their reliance on single-view representations. Most work focuses only on raw EEG signals, while others, like FBCNet [10] and IFNet [11], construct multiple filter bands to obtain various spectral views. Yet, few approaches systematically explore and fuse complementary views across time, frequency, and spatial domains. The lack of diversity hinders the model’s generalization ability, especially in cross-subject [12] and cross-headset [13] scenarios. Thus, developing principled decoding models that fuse multi-view information and encode inductive priors remains a pressing need. Multi-view prior knowledge can be integrated into data-driven MI decoding networks through data augmentations, as illustrated in Fig. 1.
Figure 1: Illustration of integrating prior knowledge into data-driven neural networks with data augmentation, e.g., temporal, spatial, and/or spectral information. Multiple views of the EEG data ensure that feature learning surpasses the limitation of designated networks. As an example, EEGNet, DCNN, and SCNN architectures rely on the intuition of CSP or filter-bank CSP for spatial variance maximization across classes (red dotted box), while not all important characteristics from time, spatial, and frequency domains of EEG data are investigated.
@@
The main contributions of this work are summarized as follows:
  1. 1.
@@
  4. 4.
    Extensive experiments on five public MI datasets across three scenarios demonstrate that MVCNet outperforms seven state-of-the-art MI decoding networks and seven data augmentation strategies, validating its effectiveness and robustness.
@@ II-A MI Decoding Networks
  1. 1.
    CNN-based models: CNNs remain the most prevalent model for EEG decoding due to their efficacy in capturing local spatial-temporal patterns. EEGNet [7] is among the most widely used models, featuring two convolutional blocks and a classification head. DCNN [8] employs four convolutional blocks with more parameters. SCNN [8], a simplified variant of DCNN, is inspired by the filter bank common spatial pattern approach. FBCNet [10] extracts spectral-spatial representations by integrating spatial convolution and temporal variance layers. Based on FBCNet, FBMSNet [14] incorporates extended convolutions for multi-scale feature extraction and introduces center loss to align features with class centers. ADFCNN [15] proposes a dual-branch CNN structure coupled with a self-attention module to enhance inter-branch feature fusion. IFNet employs two convolutional layers to extract spectral-spatial representations, similar to FBCNet.
  2. 2.
    Transformer-based models: Transformers have been introduced into EEG decoding owing to their strong capability in modeling global temporal dependencies. EEGConformer [9] adopts a sequential architecture combining SCNN with a Transformer block to capture high-level spatial-temporal patterns. Ma et al. [16] propose a comparable model integrating FBCNet with Transformer layers, similar to EEGConformer.
  3. 3.
    Mamba-based models: Mamba [17], a recently introduced alternative to Transformers, addresses the challenges of long-sequence modeling and resource inefficiencies. EEGMamba [18] utilizes a bidirectional Mamba module to effectively capture dependencies across EEG tokens. MI-Mamba [19] integrates CNN and Mamba blocks sequentially, akin to the architecture of EEGConformer.
@@ III-A Overview
This section details the proposed MVCNet. Unlike conventional single branch MI decoding models, such as CNN-based architectures (e.g., EEGNet, DCNN) or serial CNN-Transformer models (e.g., EEGConformer) that might not capture the full spectrum of EEG characteristics, MVCNet is designed to fully exploit the complementary strengths of convolutional and attention mechanisms via a parallel structure. As illustrated in Fig. 2, the input EEG trials are first augmented to generate multiple views, and then the original and augmented trials are fed into two parallel branches:
  1. 1.
@@ III-B Network Architecture
The CNN branch is designed to capture localized spatial-temporal dependencies features. It comprises a spatial convolutional layer, a temporal convolutional layer, and an average pooling operation. The spatial convolution layer adopts a depthwise separable 1D convolution equipped with $F$ spatial filters, followed with batch normalization. The temporal convolution layer applies $F$ temporal filters of size $63$, consistent with IFNet [11], and is accompanied by batch normalization and a GELU activation. Subsequently, average pooling is employed to reduce the temporal resolution and compress feature dimensionality.
The Transformer branch is configured to model temporal contextual dependencies. It consists of a lightweight self-attention encoder comprising a Transformer block, with 2 heads and 2 layers, followed by a two-layer MLP equipped with batch normalization and ReLU activation, following the design in [39]. This branch enables the network to learn long-range interactions and higher-order temporal patterns beyond local receptive fields.
@@
CNN Branch Separable Conv1D $F$ (1,)(1,)( 1 , ) $C\cdot F$ $(B,F,T)$ Spatial filtering (depthwise)
BatchNorm1D $2F$ $(B,F,T)$
Temporal Conv1D $F$ (64,)(64,)( 64 , ) $64\cdot F$ $(B,F,T)$ Temporal filtering, padding = same
BatchNorm1D $2F$ $(B,F,T)$
@@ III-C EEG Data Augmentation
  • Data flipping (Flip) [40], which flips the EEG trial of each channel in the time domain, resulting in opposite voltage values.
@@
  • Half Sample (HS) [28], which randomly selects the left brain part and the right brain part of different EEG trials, then recombines the two parts together to form a new sample.
Operations of data augmentation approaches are detailed in Table II. Visualizations of EEG trials before and after data augmentation are shown in Fig. 3.
@@ III-D Cross-View Contrasting
Given a mini-batch of $N$ trials, a total of $3N$ augmented trials are generated. For each trial, three positive pairs and $4(N-1)$ negative pairs are constructed in each branch (CNN or Transformer), resulting in $6$ positive pairs and $8(N-1)$ negative pairs in total, as illustrated in Fig. 4.
@@
We adopt the NT-Xent loss [29] as the contrastive objective, aiming to maximize the similarity between positive pairs while minimizing it for negative pairs. The contrastive distance for a sample $\mathbf{z}_{i}$ is computed as:
@@
where $v\in\{T,S,F\}$ denotes the view type, $\mathrm{sim}(\cdot,\cdot)$ denotes cosine similarity, $\tau$ is a temperature hyperparameter, and $\mathbb{I}_{[i\neq j]}$ is an indicator function that equals 1 if $i\neq j$, and 0 otherwise.
@@ III-E Cross-Model Contrasting
For each sample, one positive pair and $4(N-1)$ negative pairs are constructed per view, yielding four positive pairs and $16(N-1)$ negative pairs in total, as shown in Fig. 5.
@@
where $\mathbf{z}_{i}^{cv}$ and $\mathbf{z}_{i}^{tv}$ represent CNN and Transformer features of the same sample under view $v$.
@@ IV-B Compared Approaches
Seven MI decoding neural networks were reproduced and compared with the proposed MVCNet.
@@ IV-C Implementation
All experiments were repeated five times with random seed list $[1,2,3,4,5]$, and the average results were reported. For all datasets, models were trained for 100 epochs using the Adam optimizer with learning rate $10^{-3}$. The temperature parameter $\tau$ was set to 0.2, following [48]. $\lambda$ and $\gamma$ were fixed at 0.1 for all datasets. Batch sizes were set to 32 for baseline models, and 64 for data augmentation approaches and MVCNet, depending on the size of the training data. For MVCNet, different combinations of augmentations were applied based on the MI task: Scale, FShift, and CR were used for the BNCI2014001, Zhou2016, and Blankertz2007 datasets, while Flip, FShift, and HS were used for the BNCI2014002 and BNCI2015001 datasets.
@@ IV-D Main Results
Table IV reports the average classification accuracies (%) of proposed MVCNet and seven baseline models across five public MI datasets under three scenarios. Observe that:
  1. 1.
    CV consistently outperforms CO across all datasets and models. For example, EEGNet improves from 73.92% (CO) to 76.66% (CV), and IFNet from 81.15% to 84.90%. This trend highlights the critical role of data partitioning: CV ensures class balance and diverse representation in training, whereas CO suffers from temporal bias, resulting in limited generalization. This underscores the importance of fair and appropriate partitioning when benchmarking within-subject BCI models.
  2. 2.
    Among the seven baselines, IFNet achieves the highest average accuracy in CO (81.15%) and CV (84.90%), benefiting from frequency-aware design. EEGConformer ranks second with 78.71% and 83.91%. Under the LOSO setting, IFNet remains best (76.16%), followed by DCNN (75.46%), indicating that deeper CNNs better capture cross-subject variance.
  3. 3.
    The proposed MVCNet consistently outperforms all baselines in all scenarios and datasets. It surpasses the second best IFNet by +3.13%, +2.71%, and +2.56%, respectively. These results validate the effectiveness of our dual-branch design and multi-view contrastive framework.
  4. 4.
@@
TABLE IV: Average classification accuracies (%) of MVCNet and seven baseline models on five MI datasets in CO, CV, and LOSO scenarios. The best average performance of each dataset is marked in bold, and the second best by an underline.
Scenario Dataset EEGNet SCNN DCNN FBCNet ADFCNN EEGConformer IFNet MVCNet (Ours)
CO BNCI2014001 69.05±1.00 73.57±2.36 59.29±1.64 68.97±1.26 73.73±2.26 78.57±0.66 77.94±0.93 80.08±1.26
Zhou2016 80.13±3.35 75.03±6.15 78.03±2.37 63.33±2.29 71.42±1.95 73.87±4.51 81.70±2.08 85.47±1.16
Blankertz2007 78.79±2.78 76.71±2.09 70.00±4.12 75.93±1.31 76.07±1.41 82.29±2.62 84.00±0.57 88.14±0.76
BNCI2014002 66.07±2.76 79.07±1.96 64.07±2.70 69.50±0.95 73.00±1.95 76.21±1.46 78.29±1.68 81.42±1.82
BNCI2015001 75.58±1.69 83.71±1.34 71.08±1.82 74.92±0.97 78.75±0.62 82.63±0.54 83.83±0.90 86.29±0.64
Average 73.92 77.62 68.49 70.53 74.59 78.71 81.15 84.28
CV BNCI2014001 71.86±0.96 78.22±0.19 61.59±1.08 74.74±0.67 77.27±0.60 83.62±0.99 82.62±0.91 84.49±0.61
Zhou2016 85.30±2.12 82.90±0.90 79.68±1.13 80.01±0.50 84.95±1.51 85.67±1.52 87.78±1.01 91.72±0.27
Blankertz2007 80.21±0.77 81.11±0.72 69.83±1.51 84.14±0.67 81.10±1.75 87.43±0.33 88.23±0.50 90.81±0.64
BNCI2014002 68.90±1.37 81.24±0.50 67.17±0.31 74.53±0.68 74.63±0.75 79.30±0.33 80.21±0.52 83.34±0.55
BNCI2015001 77.04±0.72 85.58±0.42 76.89±1.05 79.57±0.27 80.69±0.40 83.53±0.71 85.67±0.56 87.68±0.24
Average 76.66 81.81 71.03 78.60 79.73 83.91 84.90 87.61
LOSO BNCI2014001 73.64±1.14 72.22±1.03 73.21±1.79 72.56±0.96 71.76±0.75 73.07±2.01 74.52±0.75 76.02±0.57
Zhou2016 83.22±1.83 82.10±0.67 83.84±1.18 82.07±0.91 82.09±1.31 82.43±1.48 86.21±0.99 86.96±1.07
Blankertz2007 71.10±0.83 70.64±0.58 72.19±0.89 76.23±1.41 70.59±1.69 74.41±1.13 73.43±1.06 79.56±1.05
BNCI2014002 72.86±0.38 70.57±1.35 74.34±0.81 71.31±0.82 72.67±0.44 72.84±1.36 73.90±0.65 75.24±0.51
BNCI2015001 71.89±0.70 69.30±0.88 73.73±0.32 67.61±0.99 71.86±0.89 70.88±0.68 72.73±1.17 75.85±0.45
Average 74.54 72.97 75.46 73.96 73.79 74.73 76.16 78.72

### v4

@@ Abstract
Electroencephalography (EEG)-based brain-computer interfaces (BCIs) enable neural interaction by decoding brain activity for external communication. Motor imagery (MI) decoding has received significant attention due to its intuitive mechanism. However, most existing models rely on single-stream architectures and overlook the multi-view nature of EEG signals, leading to limited performance and generalization. We propose a multi-view contrastive network (MVCNet), a dual-branch architecture that parallelly integrates CNN and Transformer blocks to capture both local spatial-temporal features and global temporal dependencies. To enhance the informativeness of training data, MVCNet incorporates a unified augmentation pipeline across time, frequency, and spatial domains. Two contrastive modules are further introduced: a cross-view contrastive module that enforces consistency of original and augmented views, and a cross-model contrastive module that aligns features extracted from both branches. Final representations are fused and jointly optimized by contrastive and classification losses. Experiments on five public MI datasets across three scenarios demonstrate that MVCNet consistently outperforms nine state-of-the-art MI decoding networks, highlighting its effectiveness and generalization ability. MVCNet provides a robust solution for MI decoding by integrating multi-view information and dual-branch modeling, contributing to the development of more reliable BCI systems.
@@ I Introduction
A further limitation of current MI decoding networks lies in their reliance on single-view representations. Most work focuses only on raw EEG signals, while others, like FBCNet [10] and IFNet [11], construct multiple filter bands to obtain various spectral views. Yet, few approaches systematically explore and fuse complementary views across time, frequency, and spatial domains. The lack of diversity hinders the model’s generalization ability, especially in cross-subject [12] and cross-headset [13] scenarios. Thus, developing principled decoding models that fuse multi-view information and encode inductive priors remains a pressing need. Multi-view prior knowledge can be integrated into data-driven MI decoding networks through data augmentations, as illustrated in Figure 1.
Figure 1: Illustration of integrating prior knowledge into data-driven neural networks with data augmentation, e.g., temporal, spatial, and/or spectral information. Multiple views of the EEG data ensure that feature learning surpasses the limitation of designated networks. As an example, EEGNet, DCNN, and SCNN architectures rely on the intuition of CSP or filter-bank CSP for spatial variance maximization across classes (red dotted box), while not all important characteristics from time, spatial, and frequency domains of EEG signals are investigated.
@@
The main contributions of this work are:
@@
  • Extensive experiments on five public MI datasets across three scenarios demonstrate that MVCNet outperforms nine state-of-the-art MI decoding networks and seven data augmentation strategies, validating its effectiveness and robustness.
@@ II-A MI Decoding Networks
  • CNN-based models: CNNs remain the most prevalent model for EEG decoding due to their efficacy in capturing local spatial-temporal patterns. EEGNet [7] is among the most widely used models, featuring two convolutional blocks and a classification head. DCNN [8] employs four convolutional blocks with more parameters. SCNN [8], a simplified variant of DCNN, is inspired by the filter bank common spatial pattern approach. FBCNet [10] extracts spectral-spatial representations by integrating spatial convolution and temporal variance layers. Based on FBCNet, FBMSNet [14] incorporates extended convolutions for multi-scale feature extraction and introduces center loss to align features with class centers. ADFCNN [15] proposes a dual-branch CNN structure coupled with a self-attention module to enhance inter-branch feature fusion. IFNet employs two convolutional layers to extract spectral-spatial representations, similar to FBCNet. Ma et al. [16] propose a comparable model integrating FBCNet with a temporal attention layer.
  • Transformer-based models: Transformers have been introduced into EEG decoding owing to their strong capability in modeling global temporal dependencies. EEG Conformer [9] adopts a sequential architecture combining SCNN with a Transformer block to capture high-level spatial-temporal patterns. Zhao et al. [17] utilize a convolutional module, similar to EEGNet, to extract local and spatial features, and a Transformer encoder to discern global dependencies.
  • Mamba-based models: Mamba [18], a recently introduced alternative to Transformers, addresses the challenges of long-sequence modeling and resource inefficiencies. EEGMamba [19] utilizes a bidirectional Mamba module to capture dependencies across EEG tokens. MI-Mamba [20] integrates CNN and Mamba blocks sequentially, akin to the architecture of EEG Conformer. SlimSeiz [21] combines multiple 1D CNN layers and a Mamba block to simultaneously extract temporal features at various resolutions and long-range temporal dependencies for seizure prediction.
@@
Recent studies have explored the integration of domain-specific priors to improve MI decoding performance. For example, temporal, spatial, or spectral characteristics of EEG signals have been explicitly encoded through data augmentation [22, 23], time-frequency transformations [24], frequency band modeling [11, 14], initialization with spatial filters [25], and multi-species/modality information fusion [13]. These approaches demonstrate the effectiveness of knowledge-data fusion decoding. However, most existing approaches incorporate such priors independently within specific modules or modalities. In contrast, our work unifies diverse temporal, spectral, and spatial priors into a joint learning framework via multi-view contrastive learning, thereby enabling more comprehensive representation learning and improved generalization.
@@ III-A Overview
This section details the proposed MVCNet. Unlike conventional single branch MI decoding models, such as CNN-based architectures (e.g., EEGNet, DCNN) or serial CNN-Transformer models (e.g., EEG Conformer) that might not capture the full spectrum of EEG characteristics, MVCNet is designed to fully exploit the complementary strengths of convolutional and attention mechanisms via a parallel structure. As illustrated in Figure 2, the input EEG trials are first augmented to generate multiple views, and then the original and augmented trials are fed into two parallel branches:
@@ III-B Network Architecture
The CNN branch is designed to capture localized spatial-temporal dependencies features. It comprises a spatial convolutional layer, a temporal convolutional layer, and an average pooling operation. The spatial convolution layer adopts a depthwise separable 1D convolution equipped with $F$ spatial filters, followed with batch normalization. The temporal convolution layer applies $F$ temporal filters of size 636363, consistent with IFNet [11], and is accompanied by batch normalization and a GELU activation. Subsequently, average pooling is employed to reduce the temporal resolution and compress feature dimensionality.
The Transformer branch is configured to model temporal contextual dependencies. It consists of a lightweight self-attention encoder comprising a Transformer block with 2 heads and 2 layers, followed by a two-layer MLP equipped with batch normalization and ReLU activation, following the design in [44]. A linear projection layer is then applied to normalize the output features into a common latent space, ensuring dimensional consistency with the CNN branch. This branch setup enables the network to learn long-range interactions and higher-order temporal patterns beyond local receptive fields, while facilitating effective feature alignment in the contrastive learning stage.
@@
[…]
@@ III-C EEG Data Augmentation
  1. 1.
    Data flipping (Flip) [45], which flips the EEG trial in the time domain, resulting in opposite voltage values.
  2. 2.
@@
  7. 7.
    Half Sample (HS) [33], which randomly selects the left brain part and the right brain part of different EEG trials, then recombines the two parts to form a new sample.
Operations of data augmentation approaches are detailed in Table II. Visualizations of EEG trials before and after data augmentation are shown in Figure 3.
@@ III-D Cross-View Contrasting
[…]
@@
We adopt the NT-Xent loss [34] as the contrastive objective, aiming to maximize the similarity between positive pairs while minimizing it for negative pairs. The contrastive distance for the extracted feature $\mathbf{z}_{i}$ is computed as:
@@
where $v\in\{T,S,F\}$ denotes the view type, $\mathrm{sim}(\cdot,\cdot)$ denotes cosine similarity, $N$ is the number of samples in a mini-batch, $\tau$ is a temperature hyperparameter, and $\mathbb{I}_{[i\neq j]}$ is an indicator function that equals 1 if $i\neq j$, and 0 otherwise.
@@ III-E Cross-Model Contrasting
[…]
@@
where $\mathbf{z}_{i}^{cv}$ and $\mathbf{z}_{i}^{tv}$ represent CNN and Transformer features with the same feature dimension of sample $i$ under view $v$.
@@ IV-B Compared Approaches
Nine MI decoding neural networks were reproduced and compared with the proposed MVCNet.
@@
  • SlimSeiz [21] employs multiple stacked 1D convolutional layers to extract temporal features at varying resolutions, followed by a lightweight Mamba block to model long-range dependencies. By leveraging the Mamba architecture, SlimSeiz achieves effective sequence modeling with less parameter overhead.
  • CTNet [17] introduces a sequential architecture that integrates a CNN block, similar to EEGNet, with a Transformer block. The convolutional layers capture local spatio-temporal features, while the Transformer enhances global temporal modeling.
@@ IV-C Implementation
All experiments were repeated five times with random seed list [1,2,3,4,5][1,2,3,4,5][ 1 , 2 , 3 , 4 , 5 ], and the average results were reported. For all datasets, models were trained for 100 epochs using the Adam optimizer with learning rate $10^{-3}$. The temperature parameter $\tau$ was set to 0.2, following [53]. $\lambda$ and $\gamma$ were fixed at 0.1 for all datasets. Batch sizes were set to 32 for baseline models, and 64 for data augmentation approaches and MVCNet, depending on the size of the training data. For MVCNet, different combinations of augmentations were applied based on the MI task: Scale, FShift, and CR were used for the BNCI2014001, Zhou2016, and Blankertz2007 datasets, while Flip, FShift, and HS were used for the BNCI2014002 and BNCI2015001 datasets. For all baseline models, architectural hyperparameters (e.g., kernel sizes, number of layers) were set according to their original papers to ensure fair comparison and reproducibility.
@@ IV-D Main Results
Table IV reports the average classification accuracies (%) of proposed MVCNet and nine baseline models across five public MI datasets under three scenarios. Observe that:
  • CV consistently outperforms CO across all datasets and models. For example, EEGNet improves from 73.92% (CO) to 76.66% (CV), and IFNet from 81.15% to 84.90%. This trend highlights the critical role of data partitioning: In the CO setting, temporal distribution shifts are introduced due to user fatigue, attentional drift, or electrode impedance changes, all of which can lead to degraded generalization. In contrast, each fold in the CV protocol contains a temporally contiguous segment of the recording session, ensuring realistic testing while maintaining class balance. As a result, CV yields improved generalization without violating realistic deployment assumptions, allowing the model learning a wider range of temporal patterns during training, while still preserving temporal independence between training and test sets. This underscores the importance of fair and appropriate partitioning when benchmarking within-subject EEG decoding models.
  • Among the nine baselines, IFNet achieves the highest average accuracy in CO (81.15%) and CV (84.90%), benefiting from frequency-aware design. EEG Conformer ranks second with 78.71% and 83.91%. Under the LOSO setting, IFNet remains best (76.16%), followed by DCNN (75.46%), indicating that deeper CNNs better capture cross-subject variance.
  • The proposed MVCNet consistently outperforms all baselines in all scenarios and datasets. It surpasses the second best IFNet by +3.11%, +2.71%, and +2.56%, respectively. These results validate the effectiveness of our dual-branch design and multi-view contrastive framework.
@@
TABLE IV: Average classification accuracies (%) of MVCNet and nine baseline models on five MI datasets in CO, CV, and LOSO scenarios. The best average performance of each dataset is marked in bold, and the second best by an underline.
Scenario Dataset EEGNet SCNN DCNN FBCNet ADFCNN SlimSeiz CTNet EEG Conformer IFNet MVCNet
CO BNCI2014001 69.05±1.00 73.57±2.36 59.29±1.64 68.97±1.26 73.73±2.26 68.89±2.05 73.49±2.05 78.57±0.66 77.94±0.93 83.17±0.74
Zhou2016 80.13±3.35 75.03±6.15 78.03±2.37 63.33±2.29 71.42±1.95 72.01±3.98 76.81±5.05 73.87±4.51 81.70±2.08 84.11±2.68
Blankertz2007 78.79±2.78 76.71±2.09 70.00±4.12 75.93±1.31 76.07±1.41 65.64±1.18 79.00±3.38 82.29±2.62 84.00±0.57 87.07±0.52
BNCI2014002 66.07±2.76 79.07±1.96 64.07±2.70 69.50±0.95 73.00±1.95 78.71±0.53 71.00±1.80 76.21±1.46 78.29±1.68 81.29±2.31
BNCI2015001 75.58±1.69 83.71±1.34 71.08±1.82 74.92±0.97 78.75±0.62 81.71±0.81 78.21±1.09 82.63±0.54 83.83±0.90 85.67±0.55
Average 73.92 77.62 68.49 70.53 74.59 73.39 75.70 78.71 81.15 84.26
CV BNCI2014001 71.86±0.96 78.22±0.19 61.59±1.08 74.74±0.67 77.27±0.60 75.56±1.48 75.97±0.76 83.62±0.99 82.62±0.91 84.49±0.61
Zhou2016 85.30±2.12 82.90±0.90 79.68±1.13 80.01±0.50 84.95±1.51 85.60±1.36 86.38±1.45 85.67±1.52 87.78±1.01 91.72±0.27
Blankertz2007 80.21±0.77 81.11±0.72 69.83±1.51 84.14±0.67 81.10±1.75 75.37±1.14 83.64±1.04 87.43±0.33 88.23±0.50 90.81±0.64
BNCI2014002 68.90±1.37 81.24±0.50 67.17±0.31 74.53±0.68 74.63±0.75 77.93±0.85 75.50±0.78 79.30±0.33 80.21±0.52 83.34±0.55
BNCI2015001 77.04±0.72 85.58±0.42 76.89±1.05 79.57±0.27 80.69±0.40 81.68±0.50 80.94±0.42 83.53±0.71 85.67±0.56 87.68±0.24
Average 76.66 81.81 71.03 78.60 79.73 79.23 80.49 83.91 84.90 87.61
LOSO BNCI2014001 73.64±1.14 72.22±1.03 73.21±1.79 72.56±0.96 71.76±0.75 69.62±1.32 73.40±1.22 73.07±2.01 74.52±0.75 76.02±0.57
Zhou2016 83.22±1.83 82.10±0.67 83.84±1.18 82.07±0.91 82.09±1.31 84.06±1.58 83.88±1.03 82.43±1.48 86.21±0.99 86.96±1.07
Blankertz2007 71.10±0.83 70.64±0.58 72.19±0.89 76.23±1.41 70.59±1.69 73.21±0.82 69.50±1.79 74.41±1.13 73.43±1.06 79.56±1.05
BNCI2014002 72.86±0.38 70.57±1.35 74.34±0.81 71.31±0.82 72.67±0.44 73.01±0.60 74.14±0.79 72.84±1.36 73.90±0.65 75.24±0.51
BNCI2015001 71.89±0.70 69.30±0.88 73.73±0.32 67.61±0.99 71.86±0.89 70.74±0.82 72.33±0.74 70.88±0.68 72.73±1.17 75.85±0.45
Average 74.54 72.97 75.46 73.96 73.79 74.13 74.65 74.73 76.16 78.72